Solve by any method.
(x - 10)(x - 4) = -5

Answer by stanbon(75887)

Solve by any method.
(x - 10)(x - 4) = -5
---
x^2-14x+40 = -5
---
x^2 - 14x + 45 = 0
----
Factor using the AC Method:
x^2-9x-5x+45 = 0
x(x-9)-5(x-9) = 0
(x-9)(x-5) = 0
x = 9 or x = 5
=======================
Cheers,
Stan .